1. A solution contains 1 billion 3p atoms (half life of S2p is 14.3 days) and it p-decays with a maximum energy to 32s. (a) what is the activity, in becquerels, of this solution? (b) how many atoms of phosphorus and of sulphur are in this solution in one week? (c) how long will it take for only 100 P atoms to remain in this solution?
Answer #1

Initial number of atoms No-1 billion

Half life of atoms t1/2 14.3 days

ln 2

Number of atoms at time t, N=N_0e^{-lambda t}

Activity of atoms dN

(a)

Initial activity Ao-ANİ -(14.3 . 86400) In 2 * 109 = 561 dps = 561 Bq

(b)

Number of Phosphorus atoms at time t=7,days

N- Noe t- (10)e0.712 10- 0.712bion 14.3

Number of Sulphur atoms NO-N 0.288 billion

(c)

Number of Phosphorus atoms N=100 at time t days

N=N_0e^{-lambda t}

100=(10^9)e^{-lambda t}=(10^9)e^{-rac{ln2}{14.3}*t}

e^{lambda t}=10^{7}

At In (10)

rac{ln 2}{14.3} t=lnleft(10^7 ight )=16.11809565

Time (in days) t=rac{16.11809565*14.3}{ln 2}=332.5, ext{days}